# Integrating Low-Altitude SAR Imaging into UAV Data Backhaul

### v1

@@ III-B Range Compression
Moreover, the shapes of RCM corresponding to different azimuths but identical ranges are the same, while they may differ with same azimuths at separate ranges. For example, as illustrated in Fig. 2(a), the curvatures of target 1 and target 2 are different, while those of target 2 and target 3 are the same. This indicates that applying an FFT along the azimuth dimension can align the trajectories of target 2 and target 3. In fact, all targets’ trajectories can be approximately aligned, which will be demonstrated subsequently. This motivates the azimuth FFT operation below.
@@ III-C Azimuth FFT
To that end, we may exploit the celebrated principle of stationary phase approximation (SPA) [cumming2005digital], to approach the azimuth Doppler spectrum as demonstrated in (31). However, the existing studies mainly focus on spectral approximations for continuous chirp signals, with no reports on the approximations of discrete OFDM communication signals modulated by non-constant modulus constellation symbols. For the completeness of this article, we develop its discrete form associated with the premise of SPA, as summarized below.
@@
Figure 2: Flowchart of data-aided OFDM-SAR imaging.
@@
where the last equality sign holds by substituting (35) into (40). This result demonstrates that: 1) RCM has no relationship with $y_{q}$, and 2) the quadratic coefficient in (40), i.e., $\frac{v^{2}}{2\bar{R}_{q}(K_{a}MT_{\mathrm{sym}})^{2}}$, is much less affected by $\bar{R}_{q}$, relative to $\frac{v^{2}T^{2}_{\mathrm{sym}}}{2\bar{R}_{q}}$ in (30). These findings suggest that shapes of all RCM trajectories are almost same in the RD domain, as depicted in Fig. 2(b). Consequently, we may conduct RCM correction (RCMC) by applying a Doppler-dependent range resampling (interpolation) scale, so that the target energy, which has migrated across range bins as a function of Doppler, can be realigned onto constant range cells, as depicted in Fig. 2(c). This operation significantly reduces the complexity relative to grid-wise compensation in the range-azimuth domain.
@@ III-D RCMC
If we assume perfect interpolations [^2] In practice, truncating the sinc kernel limits its accuracy. It may cause slight resolution loss, sidelobe distortion, and residual RCM errors [mao2018knowledge], especially for large fractional shifts. As we focus on the impact of signaling randomness on SAR imaging, RCM is thus assumed to be perfectly corrected., then Doppler columns contain target energy aligned to the same range index, giving rise to
@@ III-E Azimuth Compression
IV Performance Evaluation
@@ IV-A Conventional SAR Imaging Metrics
Practical SAR measurements typically deploy a corner reflector within the imaging scene and use its reconstructed response as the reference PSF [chatzitheodoridi2022cooperative]. Based on the PSF, a variety of point-target performance metrics, such as range-azimuth resolution, ISLR, NESZ, and PEL, have been introduced. By examining the imaging quality of this corner reflector, these metrics provide a standardized and comparable basis across different SAR systems and processing algorithms. In this context, the evaluation metrics of PSF generally fall into three main categories:
@@ IV-A1 Resolution‐related (mainlobe) indicators
Range and azimuth resolutions are two key metrics to distinguish two closest points in a reconstructed image. To be specific, the range resolution $\rho_{r}$ is determined by the signaling bandwidth $B_{r}$, i.e.,
$\rho_{r}=\frac{c}{2B_{r}},$ (48)
@@ IV-A2 Sidelobe indicators
The ISLR is a popular imaging metric used to quantify the total energy of the sidelobes relative to the mainlobe of PSF [stoica2009new, liu2017circulate, zhen2021multicarrier]. In the context of the signaling randomness in OFDM-SAR imaging, ISLR can be defined as
@@
However, sidelobe indicators do not take the effect of noise into account. Therefore, it cannot fairly judge the SAR imaging performance under different TF-domain filtering schemes, particularly in low SNR case.
@@ IV-A3 Radiometric quality indicators
According to [du2025prob], mismatched filtering schemes may arise an additional PEL of PSF, and amplify the noise power. To illustrate this, we may resort to the following metrics.
  • PEL: In practical SAR imaging systems, factors such as windowing, interpolation errors, truncation of the PSF kernel, motion errors, or waveform distortions can reduce the peak energy of PSF. The PEL thus serves as a measure of the system’s ability to preserve target amplitude and contrast in the reconstructed image. As we concentrate on the impact of signaling randomness on the imaging performance, we notice that both MF and RF preserve the peak energy of PSF while WF induces a peak loss, due to its scale-variant characteristic. By referring to [du2025prob], we may define PEL as
    $\mathrm{PEL}=NM\mathbb{E}\left\{\left(1-\frac{R(0,0)}{\sqrt{NM}}\right)^{2}\right\},$ (51)
@@
  • Output SNR ($\text{SNR}_{\text{out}}$): In SAR systems, the NESZ [patyuchenko2009design] is another popular metric, representing the reference RCS of a target when it equals the system noise. In other words, NESZ reflects the minimum detectable scattering, i.e., the system sensitivity, when $\mathrm{SNR_{out}=0\ dB}$. Note that $\text{SNR}_{\text{out}}$ directly represents the ratio of the mainlobe peak energy of PSF to the output noise power. Therefore, it inherently incorporates NESZ, and further accounts for the target’s scattering strength and focusing gain, thus making it a more intuitive measure of the visibility and imaging quality. $\text{SNR}_{\text{out}}$ has been defined in (18), which is equivalent to the following expression in terms of the PSF [du2025prob]:
@@ IV-B OFDM-SAR Imaging Metrics
One may exploit the MSE of the entire SAR profile as a possible imaging metric. However, it is generally not practical, since the ground-truth RCS distribution of all scatterers is unknown in real scenes, making it impossible to compute the grid/pixel-wise error between the reconstructed image and the ideal reference. Moreover, the large dynamic range of RCS values causes the MSE to be dominated by a few strong scatterers, which fails to reflect the regional image quality [^3] To address this limitation, the structural similarity index measure (SSIM) [wang2004image] is used to evaluate image quality in a perceptually meaningful way. Instead of measuring absolute intensity deviations, SSIM compares local patterns of luminance, contrast, and structure between two images within a sliding window. This normalization suppresses the dominance of strong reflections and allows for a more balanced assessment of structural fidelity across different intensity levels. Nevertheless, both MSE and SSIM are reference-dependent metrics that require a ground-truth image for comparison.. Therefore, SAR image quality is typically assessed using physical or statistical indicators that can be directly derived from the reconstructed image, such as the above mentioned point target-related metrics, which are free of reference images.
@@
For a further simplification, the first term on the right side of (53) can be derived as (54) at the top of this page. During this derivation, Dirichlet kernels are exploited [du2025prob], expressed as $\sum_{n,n^{\prime}}e^{j\frac{2\pi(n-n^{\prime})(k-k_{q})}{N}}\approx N^{2}\mathrm{sinc}^{2}(k-k_{q})$ and $\sum_{p,p^{\prime}}e^{j\frac{2\pi(p-p^{\prime})(m-m_{q})}{M}}\approx M^{2}\mathrm{sinc}^{2}(m-m_{q})$.
@@
for which a detailed proof can be found in [du2025prob, Theorem 1]. Clearly, the NMSE of the reference point target serves as a more comprehensive metric to evaluate SAR imaging quality by combining the effect of ISLR, PEL, and $\text{SNR}_{\text{out}}$. Notably, the NMSE does not account for resolution-related metrics. Consequently, one may jointly evaluate the NMSE of a reference point target and its spatial resolution within the region of interest, so as to properly characterize the imaging performance.
TABLE I: OFDM signaling parameters.
Symbol Parameter Value
${f}_{c}$ Carrier frequency 3.5 GHz
${B}_{r}$ Bandwidth 100 MHz
$\Delta f$ Subcarrier spacing 30 KHz
$T_{\rm{cp}}$} Cyclic prefix duration $8.33\ \mu s$
$T_{p}$} Transmit OFDM symbol duration $33.33\ \mu s$
$T_{a}$ Synthetic azimuth time 2 s
@@ V Simulations and Performance Analysis
We consider a UAV-SAR system in LAWNs, where the OFDM signaling parameters follow standard 5G NR configuration at the n78 frequency band, as listed in Table. I. Throughout simulations, all data symbols are i.i.d. drawn from a $256$-QAM constellation. Besides, the UAV-SAR platform, which moves at the constant speed of $v=50$ m/s along the azimuth direction, is at the height of $H_{p}=1000$ m. Notice that the data backhaul follows the standard communication transceiver chain, and thus we merely concentrate on the SAR imaging performance.
@@ V-A Flowchart of OFDM-SAR Imaging
In this subsection, we present the processing flow of the OFDM-SAR imaging procedure. To illustrate the key steps in a representative scenario, we fix the input SNR at $\mathrm{SNR_{in}}=-20$ dB and adopt the WF-based scheme as the TF filter. A detailed performance comparison among different TF-domain filtering strategies under varying $\text{SNR}_{\text{in}}$ levels is deferred to the next subsection. As shown in Fig. 3(a), three distinct RCM trajectories are observed. To map them onto a common reference scale, the SAR data are first transformed into the RD domain, where the two trajectories corresponding to identical ranges but different azimuth positions become aligned, as illustrated in Fig. 3(b). Subsequently, the RCMC operation in Fig. 3(c) further flattens these two trajectories along the azimuth dimension. Finally, azimuth compression produces the high-resolution point-target reconstruction shown in Fig. 3(d), whose focused responses closely match the true target locations. Overall, these simulation results are highly consistent with the theoretical analysis in Fig. 2, thereby validating the effectiveness of the proposed OFDM-SAR processing chain.
@@ V-B Comparison Among Different TF Filters
In the low $\text{SNR}_{\text{in}}$ regime, the filtering performance is predominantly determined by the output noise power, and MF and WF perform significantly better than RF. This is because, with RF, the element-wise division applied to non-constant modulus symbols leads to a substantial amplification of the output noise. In the medium $\text{SNR}_{\text{in}}$ regime, the sensing performance is jointly affected by the signaling randomness and the amplified noise power, while WF still provides the best focusing performance. In the high $\text{SNR}_{\text{in}}$ regime, RF and WF achieve similar sensing performance and clearly outperform MF, as the impact of randomness becomes dominant over noise and can be effectively mitigated via the reciprocal operation. Overall, WF outperforms both RF and MF across all considered $\text{SNR}_{\text{in}}$ conditions, but it requires prior knowledge of $\text{SNR}_{\text{in}}$, which introduces additional implementation overhead [^4] The transmitted power of conventional spaceborne and airborne SAR systems is typically high. In contrast, low-power, miniaturized UAV-SAR systems generally has a much lower power, especially when the communication module is multiplexed for imaging purposes. From this perspective, MF/WF may be a more suitable choice for the considered LAWNs scenario..
@@
To more precisely quantify the imaging performance among RF, MF and WF, we plot the NMSE versus $\text{SNR}_{\text{in}}$ in Fig. 5, using the phase shift keying (PSK)-modulated data symbols as the ideal sensing baseline. As expected, WF becomes equivalent to RF in the high $\text{SNR}_{\text{in}}$ regime ($\mathrm{SNR_{in}\geq 25\ dB}$) and approaches MF in the low $\text{SNR}_{\text{in}}$ regime ($\mathrm{SNR_{in}\leq-5\ dB}$). These trends are consistent with the qualitative observations in Fig. 4.
@@ V-C Comparison Among Pilot-only and Data-aided Schemes
We now demonstrate the potential of exploiting data-aided imaging and illustrate its superiority over the pilot-only scheme. Unless otherwise specified, MF is employed under $\text{SNR}_{\text{in}}$=5 dB in this subsection. As a benchmark, the pilot-only scheme relies on SRS, which is mapped onto the OFDM subcarriers using a comb-pattern, resulting in a sparse, comb-shaped frequency spectrum. Referring to the 5G NR uplink frame structure [etsi38138], the SRS occupies $24$ resource blocks, where each block contains $12$ consecutive subcarriers. Therefore, the total occupied bandwidth of the SRS is $24\times 12\times\Delta f=8.64$ MHz, which is less than one-tenth of $B_{r}$. In the simulation, the SRS subcarrier indices range from $1667$ to $1954$. It is worth noting that the comb-pattern is repeated with a frequency spacing of 4 subcarriers. Additionally, SRS can be transmitted with various periodicities. Below we consider two SRS configurations, i.e., a long periodicity of $20$ slots and a short periodicity of $2$ slots, where each slot contains $14$ symbols.
@@
(a) Pilot-only: periodicity of 20 slots
(b) Pilot-only: periodicity of 2 slots
(c) Data-aided imaging
(d) Range and azimuth profiles
@@
As shown in Fig. 6(a) and Fig. 6(d), the pilot-only imaging scheme suffers from severe defocusing: multiple peaks appear in the entire range-azimuth image, leading to significant distortion in the response of reference point target. We attribute this behavior to two reasons. First, in the azimuth domain, the PRF of the SRS is $\frac{1}{20\times 14\times T_{\mathrm{sym}}}\approx 85.7\ \mathrm{Hz}<2B_{a}$, which results in additional azimuth peaks due to spectral aliasing caused by undersampling. However, for the SRS configuration of $2$ slots’ periodicity, the spectral aliasing in the azimuth can be eliminated due to oversampling, as illustrated in Fig. 6(b) and Fig. 6(d). Second, periodic peaks along the range dimension occur because of the discontinuous comb-structure across subcarriers [liu2016space], where the adjacent peak interval is $\frac{1}{4\Delta f}=8.33\ \mu s$, corresponding to a round-trip range of 1250 m, as observed in Fig. 6(a), Fig. 6(b) and Fig. 6(d). In contrast, the data-aided scheme yields a well-focused image in Fig. 6(c), which significantly outperforms the pilot-only benchmark. The involvement of more data symbols enhances the energy accumulation, thereby providing a higher SNR gain. Additionally, Fig. 6(d) confirms that the azimuth resolution is preserved due to the fixed $T_{a}$, while the range resolution can be significantly facilitated with data-aided scheme, as it occupies the entire $B_{r}$.
@@
In Fig. 8, we further reconstruct the SAR image of an extended scene composed of several isolated point scatterers. In this simulation, we employ a real binary grayscale scene illustrated in Fig. 8(a), which is used as the reference image for SAR data reconstruction. Specifically, the imaging performance with pilot-only schemes is clearly unsatisfactory, as depicted in Fig. 8(b) and Fig. 8(c). This is because SRS occupies only a small fraction of the available bandwidth, leading to limited achievable range resolution. Additionally, the fewer symbols involved leads to a modest SNR gain. Consequently, the reconstructed scene suffers from noticeable blurring and distortion. In contrast, the data-aided scheme fully exploits the entire bandwidth and a denser set of symbols over the synthetic aperture, which jointly enhance both the effective range resolution and the SNR, without additional pseudo peaks. As a result, the point scatterers are much better focused and more faithfully reconstructed, demonstrating a pronounced performance improvement over the pilot-only imaging scheme in practical SAR scenes.
(a) Reference image

### v2

@@ III-B Range Compression
Moreover, the shapes of RCM corresponding to different azimuths but identical ranges are the same, while they may differ with same azimuths at separate ranges. In fact, all targets’ trajectories can be approximately aligned, which will be demonstrated subsequently. This motivates the azimuth FFT operation below.
@@ III-C Azimuth FFT
To that end, we may exploit the celebrated principle of stationary phase approximation (SPA) [5], to approach the azimuth Doppler spectrum as demonstrated in (31). However, the existing studies mainly focus on spectral approximations for continuous chirp signals, with no reports on the approximations of discrete OFDM communication signals modulated by non-constant modulus constellation symbols. For the completeness of this article, we develop its discrete form associated with the premise of SPA, as summarized below.
@@
where the last equality sign holds by substituting (35) into (40). This result demonstrates that: 1) RCM has no relationship with $y_{q}$, and 2) the quadratic coefficient in (40), i.e., $\frac{v^{2}}{2\bar{R}_{q}(K_{a}MT_{\mathrm{sym}})^{2}}$, is much less affected by $\bar{R}_{q}$, relative to $\frac{v^{2}T^{2}_{\mathrm{sym}}}{2\bar{R}_{q}}$ in (30). These findings suggest that shapes of all RCM trajectories are almost same in the RD domain. Consequently, we may conduct RCM correction (RCMC) by applying a Doppler-dependent range resampling (interpolation) scale, so that the target energy, which has migrated across range bins as a function of Doppler, can be realigned onto constant range cells. This operation significantly reduces the complexity relative to grid-wise compensation in the range-azimuth domain.
@@ III-D RCMC
If we assume perfect interpolations [^4] In practice, truncating the sinc kernel limits its accuracy. It may cause slight resolution loss, sidelobe distortion, and residual RCM errors [26], especially for large fractional shifts. As we focus on the impact of signaling randomness on SAR imaging, RCM is thus assumed to be perfectly corrected., then Doppler columns contain target energy aligned to the same range index, giving rise to
@@ III-E Azimuth Compression
IV Mapping Mechanism Between Data Randomness and SAR Imaging
@@ IV-A Conventional SAR Imaging Metrics
Practical SAR measurements typically deploy a corner reflector within the imaging scene and use its reconstructed response as the reference PSF [3]. Based on the PSF, a variety of point-target performance metrics, such as range-azimuth resolution, integrated sidelobe ratio (ISLR) [33], noise-equivalent sigma zero (NESZ) [28], and peak energy loss (PEL) [9], have been introduced. By examining the imaging quality of this corner reflector, these metrics provide a standardized and comparable basis across different SAR systems and processing algorithms. In this context, the evaluation metrics of PSF generally fall into three main categories:
@@ IV-A1 Resolution‐related (mainlobe) indicators
Range and azimuth resolutions are two key metrics to distinguish two closest points in a reconstructed image. Specifically, the range resolution $\rho_{r}$ depends on the signaling bandwidth $B_{r}$, i.e.,
$\rho_{r}={c}/(2B_{r}),$ (48)
@@ IV-A2 Sidelobe indicators
The ISLR is a popular imaging metric used to quantify the total energy of the sidelobes relative to the mainlobe of PSF [33, 23, 11]. In the context of the signaling randomness in OFDM-SAR imaging, ISLR can be defined as
@@
Overall, sidelobe indicators are primarily designed to characterize the PSF shape under well-focused conditions, and thus do not explicitly account for noise effects. In conventional SAR systems, noise-related performance is typically assessed using separate radiometric metrics such as NESZ or output SNR. Therefore, rather than indicating a limitation of ISLR, this reflects its specific role in evaluating sidelobe behavior.
@@ IV-A3 Radiometric quality indicators
According to [9], mismatched filtering schemes may arise an additional PEL of PSF, and amplify the noise power. To illustrate this, we may resort to the following metrics.
$\bullet$ PEL: In practical SAR imaging systems, factors such as windowing, interpolation errors, truncation of the PSF kernel, motion errors, or waveform distortions can reduce the peak energy of PSF. The PEL thus serves as a measure of the system’s ability to preserve target amplitude and contrast in the reconstructed image. As we concentrate on the impact of signaling randomness on the imaging performance, we notice that both MF and RF preserve the peak energy of PSF while WF induces a peak loss, due to its scale-variant characteristic. By referring to [9], we may define PEL as
$\mathrm{PEL}=NM\mathbb{E}\left\{\left(1-{R(0,0)}/{\sqrt{NM}}\right)^{2}\right\},$ (51)
@@
$\bullet$ Output SNR: In SAR systems, the NESZ [28] represents the RCS of a reference target when it equals the system noise. In other words, NESZ reflects the minimum detectable scattering, i.e., the system sensitivity, when $\mathrm{SNR_{out}=0\ dB}$. Note that $\text{SNR}_{\text{out}}$ directly represents the ratio of the mainlobe peak energy of PSF to the output noise power. Therefore, it inherently incorporates NESZ, and further accounts for the target’s scattering strength and focusing gain, thus making it a more intuitive measure of the visibility and imaging quality. $\text{SNR}_{\text{out}}$ has been defined in (18), which is equivalent to the following expression in terms of the PSF [9]:
@@ IV-B OFDM-SAR Imaging Metrics
In scenarios where both sidelobe structure and noise are affected by the modulation randomness and TF-filtering schemes, it is desirable to adopt a unified metric that jointly captures these effects. To this end, one may exploit the MSE of the entire SAR profile as a possible imaging metric. However, it is generally not practical, since the ground-truth RCS distribution of all scatterers is unknown in real scenes, making it impossible to compute the grid/pixel-wise error between the reconstructed image and the ideal reference. Moreover, the large dynamic range of RCS values causes the MSE to be dominated by a few strong scatterers, which fails to reflect the regional image quality [^5] To address this limitation, the structural similarity index measure (SSIM) [39] is used to evaluate image quality in a perceptually meaningful way. Instead of measuring absolute intensity deviations, SSIM compares local patterns of luminance, contrast, and structure between two images within a sliding window. This normalization suppresses the dominance of strong reflections and allows for a more balanced assessment of structural fidelity across different intensity levels. Nevertheless, both MSE and SSIM are reference-dependent metrics that require a ground-truth image for comparison.. Therefore, SAR image quality is typically assessed using physical or statistical indicators that can be directly derived from the reconstructed image, such as the above mentioned point target-related metrics, which are free of reference images.
@@
For a further simplification, the first term on the right side of (53) can be derived as (54) at the top of this page. During this derivation, Dirichlet kernels are exploited [9], expressed as $\sum_{n,n^{\prime}}e^{j\frac{2\pi(n-n^{\prime})(k-k_{q})}{N}}\approx N^{2}\mathrm{sinc}^{2}(k-k_{q})$ and $\sum_{p,p^{\prime}}e^{j\frac{2\pi(p-p^{\prime})(m-m_{q})}{M}}\approx M^{2}\mathrm{sinc}^{2}(m-m_{q})$.
@@
for which a detailed proof can be found in [9, Theorem 1]. In the considered OFDM-SAR framework, NMSE provides a unified measure that jointly reflects the impact of sidelobe structure, peak distortion, and noise, which are otherwise characterized separately by ISLR, PEL, and SNR-related metrics.
It is worth noting that NMSE is not intended to replace conventional SAR performance metrics such as resolution, ISLR, or radiometric indicators. Instead, it serves as a complementary metric that is particularly useful in scenarios where a reference point target is available, such as calibration experiments or controlled simulations. For example, to realize trade-off between sensing and communications via waveform design, NMSE can be directly optimized through probabilistic constellation shaping (PCS) [8], offering a much simpler alternative to optimizing multiple conventional metrics such as ISLR, PEL, and $\mathrm{SNR_{out}}$. We refer to [9] for more details. Therefore, it is suitable for comparative performance evaluation of different waveform and filtering strategies, and should be interpreted together with conventional SAR metrics for a comprehensive assessment of imaging quality.
@@ V Simulations and Performance Analysis
We consider a UAV-SAR system in LAWNs, where the OFDM signaling parameters follow standard 5G NR configuration at the n78 frequency band, where the carrier frequency is ${f}_{c}=3.5$ GHz, the bandwidth is $B_{r}=100$ MHz, the subcarrier spacing is $\Delta f=30$ KHz, the cyclic prefix duration is ${T_{\rm{cp}}}=8.33\ \mu s$, the transmit OFDM symbol duration is ${T_{p}}=33.33\ \mu s$, and the synthetic azimuth time $T_{a}=2$ s. Throughout simulations, all data symbols are i.i.d. drawn from a $256$-QAM constellation. Besides, the UAV-SAR platform, which moves at the constant speed of $v=50$ m/s along the azimuth direction, is at the height of $H_{p}=1000$ m. Notice that the data backhaul follows the standard communication transceiver chain, and thus we merely concentrate on the SAR imaging performance.
@@ V-A Flowchart of OFDM-SAR Imaging
In this subsection, we present the processing flow of the OFDM-SAR imaging procedure. To illustrate the key steps in a representative scenario, we fix the input SNR at $\mathrm{SNR_{in}}=-20$ dB and adopt the WF-based scheme as the TF filter. A detailed performance comparison among different TF-domain filtering strategies under varying $\text{SNR}_{\text{in}}$ levels is deferred to the next subsection. As shown in Fig. 3(a), three RCM trajectories exist. To map them onto a common reference scale, the SAR data are first transformed into the RD domain, where the two trajectories corresponding to identical ranges but different azimuth positions become aligned, as illustrated in Fig. 3(b). Subsequently, the RCMC in Fig. 3(c) further flattens these two trajectories along the azimuth dimension. Finally, the azimuth compression produces the high-resolution point-target reconstruction shown in Fig. 3(d), whose focused responses closely match the true target locations.
@@ V-B Comparison Among Different TF Filters
In the low $\text{SNR}_{\text{in}}$ regime, the filtering performance is predominantly determined by the output noise power, and MF and WF perform significantly better than RF. This is because, with RF, the element-wise division applied to non-constant modulus symbols leads to a substantial amplification of the output noise. In the medium $\text{SNR}_{\text{in}}$ regime, the sensing performance is jointly affected by the signaling randomness and the amplified noise power, while WF still provides the best focusing performance. In the high $\text{SNR}_{\text{in}}$ regime, RF and WF achieve similar sensing performance and clearly outperform MF, as the impact of randomness becomes dominant over noise and can be effectively mitigated via the reciprocal operation. Overall, WF outperforms both RF and MF across all considered $\text{SNR}_{\text{in}}$ conditions [^6] The transmitted power of conventional spaceborne and airborne SAR systems is typically high. In contrast, low-power, miniaturized UAV-SAR systems generally has a much lower power, especially when the communication module is multiplexed for imaging purposes. From this perspective, MF/WF may be a more suitable choice for the considered LAWNs scenario..
@@
To more precisely quantify the imaging performance among RF, MF and WF, we plot the NMSE versus $\text{SNR}_{\text{in}}$ in Fig. 5, using the phase shift keying (PSK)-modulated data symbols as the sensing baseline. As expected, WF becomes equivalent to RF in high $\text{SNR}_{\text{in}}$ regime ($\mathrm{SNR_{in}\geq 25\ dB}$) and approaches MF in low $\text{SNR}_{\text{in}}$ regime ($\mathrm{SNR_{in}\leq-5\ dB}$). These trends are consistent with the qualitative observations in Fig. 4.
@@ V-C Comparison Among Pilot-only and Data-aided Schemes
We now explore the potential of data-aided imaging and illustrate its superiority over the pilot-only scheme. Unless otherwise specified, MF is employed under $\text{SNR}_{\text{in}}$=5 dB in this subsection. As a benchmark, the pilot-only scheme relies on SRS, which is mapped onto the OFDM subcarriers using a comb-pattern, resulting in a sparse, comb-shaped frequency spectrum. Referring to the 5G NR uplink frame structure [12], the SRS occupies $24$ resource blocks, where each block contains $12$ consecutive subcarriers. Therefore, the total occupied bandwidth of the SRS is $24\times 12\times\Delta f=8.64$ MHz, which is less than one-tenth of $B_{r}$. In the simulation, the SRS subcarrier indices range from $1667$ to $1954$. It is worth noting that the comb-pattern is repeated with a frequency spacing of 4 subcarriers. Additionally, SRS can be transmitted with various periodicities. Below we consider two SRS configurations, i.e., a long periodicity of $20$ slots and a short periodicity of $2$ slots, where each slot contains $14$ symbols.
@@
As shown in Fig. 6, the pilot-only imaging scheme suffers from severe defocusing: multiple peaks appear in the entire range-azimuth image, leading to significant distortion in the response of reference point target. We attribute this behavior to two reasons. First, in the azimuth domain, the PRF of the SRS is $\frac{1}{20\times 14\times T_{\mathrm{sym}}}\approx 85.7\ \mathrm{Hz}<2B_{a}$, which results in additional azimuth peaks due to spectral aliasing caused by undersampling. It is worth noting that the aliasing peaks (with the SRS configuration of $20$ slots’ periodicity) are substantially suppressed relative to the target peak, owing to the modulation of the azimuth beampattern, which generally exhibits a $\text{sinc}^{2}$ shape [5]. However, for the SRS configuration of $2$ slots’ periodicity, the spectral aliasing in the azimuth can be eliminated due to oversampling. Second, periodic peaks along the range dimension occur because of the discontinuous comb-structure across subcarriers [22], where the adjacent peak interval is $\frac{1}{4\Delta f}=8.33\ \mu s$, corresponding to a round-trip range of 1250 m. In contrast, the data-aided scheme yields a well-focused image, which significantly outperforms the pilot-only benchmark. The involvement of more data symbols enhances the energy accumulation, thereby providing a higher SNR gain. Additionally, Fig. 6 confirms the identical azimuth resolution of three schemes due to the fixed $T_{a}$, while the range resolution can be significantly facilitated with data-aided scheme, as it occupies the entire $B_{r}$.
@@
In Fig. 8, we further reconstruct the SAR image of an extended scene, where a real SAR image [38] in Fig. 8(a) as the reference image for SAR data reconstruction. Evidently, the imaging performance with pilot-only schemes is unsatisfactory, as depicted in Fig. 8(b) and Fig. 8(c). This is because SRS occupies only a small fraction of the available bandwidth, leading to limited achievable range resolution. Additionally, the fewer symbols involved leads to a modest SNR gain. Consequently, the reconstructed scene suffers from noticeable blurring and distortion. In contrast, the data-aided scheme fully exploits the entire bandwidth and a denser set of symbols over the synthetic aperture, which jointly enhance both the effective range resolution and the SNR, without additional pseudo peaks. As a result, the point scatterers are much better focused and more faithfully reconstructed, demonstrating a pronounced performance improvement over the pilot-only imaging scheme in practical SAR scenes.
(a) Reference SAR image [38]